OpenStudy (anonymous):

Solve -4(x + 10) - 6 = -3(x - 2).

OpenStudy (anonymous):

first simplify to -4x-46=-3x+6

OpenStudy (anonymous):

dont give direct answer @eninone

OpenStudy (baby_bear69):

First, distribute and simplify. so -4x-40-6= -3x=6

OpenStudy (baby_bear69):

Ah you beat me to the send button @DaBest21

OpenStudy (anonymous):

thanks guy but i already got it 

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

then get the variables on one side

OpenStudy (anonymous):

@DaBest21 Ok, dear

OpenStudy (anonymous):

-4x-46=-3x+6 +4x +4x -46=x+6 -6 -6 -52=x

OpenStudy (anonymous):

to check plug -52 in for x

OpenStudy (anonymous):

Okay thank you so much :) @DaBest21

OpenStudy (anonymous):

np

OpenStudy (anonymous):

\[-4(x+10)-6=-3(x-2)\\ -4x-40-6=-3x+6\\-4x+3x=40+6+6\\-x=52\\x=-52 \ Ans.\]

OpenStudy (anonymous):

Don't forget to medal and fan if you already haven't. Plus mention my name for help @Divamygirl